OpenStudy (anonymous):

Derivative of population formula. The formula is that of a butterfly population. It is y=6000/(1+46(0.6)^t)

OpenStudy (anonymous):

I read the answer in the book but they don't explain the steps so I don't understand how they got the answer.

OpenStudy (anonymous):

is it \[\frac{6000}{1+46(.6)^t}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I changed its form for product rule applciation

OpenStudy (anonymous):

ok the derivative of \(\frac{1}{f}\) is \(-\frac{f'}{f^2}\)

OpenStudy (anonymous):

But this requires us to use chain rule as well and no matter how many times I do it mself I get the wrong answer

OpenStudy (anonymous):

nah, product is too confusing use \[-\frac{f'}{f^2}\]

OpenStudy (anonymous):

I amde it 6000(1+49(0.6)^t)^-1

OpenStudy (anonymous):

I have never seen that what is that rule called

OpenStudy (anonymous):

\[f(x)=1+46(.6)^t\] \[f'(x)=46\ln(.6)(.6^t)\]

OpenStudy (anonymous):

In the book they sue the product rule and the chain rule so ideally i wanna learn how to do it the way they did it

OpenStudy (anonymous):

use*

OpenStudy (anonymous):

you are making too much work for yourself

OpenStudy (anonymous):

which rule?

OpenStudy (anonymous):

f′(x)=46ln(.6)(.6t) <---- what rule is that! Where did you get the ln!

OpenStudy (anonymous):

the derivative of \(b^x\) is \(\ln(b)\times b^x\) always

OpenStudy (anonymous):

Can you explain to me in steps how you would do the whole things using product rule and chian rule step by step Ill follow on paper. Sorry for the hassle but I really need to internalize this process

OpenStudy (anonymous):

yeah hold on one minute

OpenStudy (anonymous):

I SUPER appreciate this btw

OpenStudy (anonymous):

ok first of all we are not going to use the product rule or the quotient rule

OpenStudy (anonymous):

even tho the book has their answer that way?

OpenStudy (anonymous):

if i had \[\frac{x-1}{x+1}\] i would use the quotient rule, because there is a variable in the numerator

OpenStudy (anonymous):

but you have no variable in the numerator, just a number (constant)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

if you want the derivative of \[\frac{c}{f}\] where \(c\) is a constant, it is \[-\frac{cf'}{f^2}\]

OpenStudy (anonymous):

the quotient rule will give you the same thing, because the derivative of the numerator is zero if it is a constant

OpenStudy (anonymous):

I guess the constant is kind of throwing me off

OpenStudy (anonymous):

so rather than using \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] if \(f\) is a number \(c\) , you just get \[-\frac{cg'}{g^2}\]

OpenStudy (anonymous):

c=-f right

OpenStudy (anonymous):

another way to look at this is as a chain rule problem \[\frac{d}{dx}\frac{1}{x}=-\frac{1}{x^2}\] and so \[\frac{d}{dx}\frac{1}{f(x)}=-\frac{1}{f^2(x)}\times f'(x)=-\frac{f'(x)}{f^2(x)}\]

OpenStudy (anonymous):

in order to do your problem, the first thing you do is put parentheses around the denominator and square it

OpenStudy (anonymous):

i don't mean actually square it, i just mean write it

OpenStudy (anonymous):

1+49^2(0.6)^2t

OpenStudy (anonymous):

\[f(t)=\frac{6000}{1+46(.6)^t}\] \[f'(t)=\frac{-6000f'(t)}{(1+46\times .6^t)^2}\]

OpenStudy (anonymous):

oh sorry ok i got it

OpenStudy (anonymous):

ok and then

OpenStudy (anonymous):

bad algebra there in any case you can't square like that....

OpenStudy (anonymous):

ya i screwed up i noticed

OpenStudy (anonymous):

the only think that is missing is \(f'(t)\)

OpenStudy (anonymous):

by which i mean the derivative of the denominator \[1+46\times .6^t\]

OpenStudy (anonymous):

Bleh your explanation is great but i keep blanking out. I'm gonna review previous units I just took a long break and forgot a bunch of rules.

OpenStudy (anonymous):

Thanks for the help, I just think I gotta drill derivative rules again

OpenStudy (anonymous):

this is what you need to know to get that derivative the derivative of 1 is 0 the derivative of \(b^x\) is \(b^x\ln(b)\)

OpenStudy (anonymous):

so the derivative of your denominator is \[46\ln(.6)\times .6^t\]

OpenStudy (anonymous):

ya i understood that part thanks

OpenStudy (anonymous):

ok then you are done answer \[\frac{-6000\times 46\ln(.6)\times .6^t}{(1+46\times .6^t)^2}\]

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thx buddy